OpenStudy (anonymous):

Which of the following shows the right side of the equation after applying a reasonable first step to solve two-thirds minus h divided by two equals five-eighths?

OpenStudy (anonymous):

(2/3 - h)/2 = 5/8 ? or 2/3 - (h/2) = 5/8 ? which one is your problem ?